Determine 3 distinct and positive numbers a,b,c forming a geometric sequence verifying ac=49 and 7a+6b-c=0
Answer by richwmiller(17219)

Tn = t * r^(n - 1)
a = t,
b = t * r,
c = t * r^2,
ac=49,
7a+6b-c=0
a = 1, b = 7, c = 49, r = 7, t = 1